# RadarCam-Depth: Radar-Camera Fusion for Depth Estimation with Learned Metric Scale

### 09 Jan 2024 v1

@@ Abstract
We present a novel approach for metric dense depth estimation based on the fusion of a single-view image and a sparse, noisy Radar point cloud. The direct fusion of heterogeneous Radar and image data, or their encodings, tends to yield dense depth maps with significant artifacts, blurred boundaries, and suboptimal accuracy. To circumvent this issue, we learn to augment versatile and robust monocular depth prediction with the dense metric scale induced from sparse and noisy Radar data. We propose a Radar-Camera framework for highly accurate and fine-detailed dense depth estimation with four stages, including monocular depth prediction, global scale alignment of monocular depth with sparse Radar points, quasi-dense scale estimation through learning the association between Radar points and image patches, and local scale refinement of dense depth using a scale map learner. Our proposed method significantly outperforms the state-of-the-art Radar-Camera depth estimation methods by reducing the mean absolute error (MAE) of depth estimation by 25.6% and 40.2% on the challenging nuScenes dataset and our self-collected ZJU-4DRadarCam dataset, respectively.
@@ I Introduction
Perceiving the environment is critically important for autonomous driving, where accurate depth estimation is fundamental for dense reconstruction, 3D detection, and motion planning. Cameras and range sensors have been widely used for perceiving dense depth. Learned monocular depth (mono-depth) estimation methods based on CNN networks [1, 2, 3, 4, 5] have been prevalent in recent years due to their versatile applicability and plausible accuracy. They benefit from the solid contextual priors from extensive training on diverse datasets. While mono-depth networks excel in estimating up-to-scale depth, their performance degrades when predicting metric depth. This limit arises from the inherent challenge of capturing scale with single-view cameras and the difficulty of learning the diverse scale in complex scenarios.
@@
Figure 1: Top: 3D visualization of the metric depth estimation of our proposed RadarCam-Depth; Middle: Our metric depth estimation overlaid on corresponding error map; Bottom: Mono-depth aligned to Radar points with a global scale factor. Our depth estimation exhibits exceptional metric accuracy and recovers fine details.
@@
Figure 2: The overall framework of our proposed RadarCam-Depth, comprised with four stages: monocular depth prediction, global alignment of mono-depth with sparse Radar depth, learned quasi-dense scale estimation, and scale map learner for refining local scale of dense depth estimation. $\mathbf{d}$ and $\mathbf{s}$ denotes the depth and scale, while $\mathbf{z}=1/\mathbf{d}$ is the inverse depth.
The primary contributions of this work are as follows: (i) We introduce the first approach that enhances the highly generalizable, scaleless mono-depth prediction with the dense metric scale intricately inferred from the noisy and sparse Radar data. (ii) We present a novel metric dense depth estimation framework that effectively fuses heterogeneous Radar and camera data. Our framework comprises four stages: monocular depth prediction, global scale alignment of the monocular depth, Radar-Camera quasi-dense scale estimation, and scale map learner for refining the quasi-dense scale locally. (iii) The proposed method is extensively tested on the nuScenes benchmark and our self-collected ZJU-4DRadarCam dataset. It outperforms the state-of-the-art (SOTA) techniques, substantially enhancing Radar-Camera dense depth estimation with high metric accuracy and solid generalizability. (iv) To address the absence of a 4D Radar-based depth estimation dataset, we will release our high-quality ZJU-4DRadarCam dataset, including raw 4D Radar data, RGB images, and meticulously generated ground-truth depth maps from LiDAR measurements. Additionally, our code will be open-sourced to fertilize further research.
@@ II-A Monocular Depth Estimation
Monocular depth estimation is a challenging task due to the inherent scale ambiguity. Despite its simplicity and the availability of raw data, researchers have made significant efforts to address this issue by integrating it with optical flow [14], uncertainty estimation [15], semantic segmentation [16], instance segmentation [17] and visual odometry [18]. Although some previous studies [19, 20, 4, 3] have achieved promising results in affine-invariant scaleless depth estimation across diverse datasets, recovering the metric scale still poses a significant challenge. Some existing methods rely on inertial data to provide scale. To enhance the generalization, VI-SLAM [21] warps the input image to match the orientation prevailing in the training dataset. Zuo et al.’s CodeVIO [22] proposes a tightly coupled VIO system with optimizable learned dense depth. It jointly estimates VIO poses and optimizes the predicted and encoded dense depth of certain keyframes efficiently. Similarly, Xie [23] utilizes a flow-to-depth layer to refine camera poses and generate depth proposals. They solve a multi-frame triangulation problem to enhance the estimation accuracy. Recently, Wofk [5] introduced a framework for metric dense depth estimation from the VIO sparse depth and monocular depth prediction, which inspires our work. They first globally align the scaleless mono-depth with the metric VIO sparse depth, and then learn to refine the dense scale of the globally aligned mono-depth.
@@ II-B Depth Estimation from Radar-Camera Fusion
Fusion of Radar and camera data for metric depth estimation is an area of active research. Lin et al. [8] introduced a two-stage CNN-based pipeline that combines Radar and camera inputs to denoise Radar signals and estimate dense depth. Long et al. [9] proposed a Radar-2-Pixel (R2P) network that utilizes radial Doppler velocity and induced optical flow from images to associate Radar points with corresponding pixel regions, enabling the synthesis of full-velocity information. They also achieved image-guided depth completion using Radar and video data [10]. Another approach, DORN [11] proposed by Lo et al, extends Radar points in the elevation dimension and applies deep ordinal regression network-based [24] feature fusion. Unlike other methods, R4dyn [13] creatively incorporates Radar as a weakly supervised signal into a self-supervised framework and employs Radar as an additional input to enhance the robustness. However, their method primarily focuses on vehicle targets and does not fully correlate all Radar points with a larger image area, resulting in lower depth accuracy. However, the above methods require multi-frame information to overcome the sparsity of Radar data. In contrast, Singh et al. [12] presented a fusion method that relies solely on a single image frame and Radar point cloud. Their first-stage network infers the confidence scores of associating a Radar point to image patches, which leads to a semi-dense depth map after association. They further employ a gated fusion network to control the fusion of multi-modal Radar-Camera data, and predict the final dense depth.
@@ III METHODOLOGY
Our goal is to recover the dense depth $\mathbf{\hat{d}}\in\mathbb{R}^{H_{0}\times W_{0}}_{+}$ from a pair of RGB image $\mathbf{I}\in\mathbb{R}^{3\times H_{0}\times W_{0}}$ and Radar point cloud $\mathbf{P}=\left\{\mathbf{p}_{i}|\mathbf{p}_{i}\in\mathbb{R}^{3},i=0,1,2,\cdots,k-1\right\}$ in the image coordinate. Either 3D or 4D Radar usually has a small field of view with ambiguous sparse data deteriorated by intensive noises. To enable the cross-modal fusion, it is straightforward to project Radar points onto the image plane, generating Radar depth. However, direct fusion of the inherently ambiguous and sparse Radar depth with images, achieved by concatenating their encodings or raw data, can confuse the learning pipeline [25, 26, 27, 12], which results in aliasing and other undesirable artifacts in the estimated depth. In this paper, we propose to get the scaleless dense depth with existing versatile monocular depth prediction networks, then learn to augment scaleless dense depth with accurate dense metric scales from Radar data.
@@ III-A Monocular Depth Prediction
Figure 3: Mono-depth predictions by MiDaS v3.1 [3] and DPT-Hybrid [20]. Notably, MiDaS exhibits the ability to differentiate the sky.
@@ III-B Global Alignment
We align the scaleless monocular depth prediction $\mathbf{\hat{d}}_{m}$ with the Radar depth originating from projecting raw Radar points $\mathbf{P}$, by a global scaling factor $\hat{s}_{g}$ and optional offset $\hat{t}_{g}$. The global aligned metric monocular depth is calculated by $\mathbf{\hat{d}}_{ga}=\hat{s}_{g}\cdot\mathbf{\hat{d}}_{m}+\hat{t}_{g}$. The globally aligned mono-depth is fed into the subsequent scale map learner (SML). There are many options for performing this global alignment between the projected Radar depth and mono-depth prediction, including: (i) Var: A varying $\hat{s}_{g}$ for individual frame of mono-depth, calculated via root-finding algorithms [29, 30]. (ii) Const: A constant $\hat{s}_{g}$ for all frames of mono-depth prediction, deemed as the mean of scale estimates on the entire training samples. (iii) LS: $\hat{s}_{g}$ and $\hat{t}_{g}$ for individual frames, computed with linear least-squares optimization [19]. (iv) RANSAC: $\hat{s}_{g}$ and $\hat{t}_{g}$ for individual frames, computed with linear least-squares while incorporating RANSAC outlier rejection of the Radar depth. We randomly sample 5 Radar points with valid values, estimate $\hat{s}_{g}$ and $\hat{t}_{g}$ with the sampled Radar depth, and adopt the first pair of $\hat{s}_{g}$ and $\hat{t}_{g}$ that yields an inlier ratio over $90\%$. Suppose the discrepancy between Radar depth and aligned mono-depth at the corresponding pixel location is under $6$m (or the inverse depth discrepancy is under $0.015$). In that case, it is a pair of inlier associations.
@@
Figure 4: (a) Top: nuScenes dataset [31] with LiDAR depth $\mathbf{d}_{gt}$ and accumulated LiDAR depth $\mathbf{d}_{acc}$, depth from 3D Radar point cloud $\mathbf{P}$, interpolated LiDAR $\mathbf{d}_{int}$ shown clockwise. The misalignment between LiDAR points and image pixels on this dataset is highlighted with red boxes. Depth from 3D Radar point cloud is very sparse and non-uniformly distributed. (b) Our ZJU-4DRadarCam dataset with LiDAR depth $\mathbf{d}_{gt}$, interpolated LiDAR depth $\mathbf{d}_{int}$, and depth from 4D Radar point cloud $\mathbf{P}$ shown from top to bottom. Compared to nuScenes, ZJU-4DRadarCam dataset offers more accurate and denser LiDAR depth, as well as denser 4D Radar depth.
@@ III-C2 Confidence of Cross-Modal Associations
$\hat{\mathbf{d}}_{q}(u,v)=\begin{cases}d(p_{\mu}),\ \ \ \ \text{if}\ \mathbf{\hat{y}}_{\mu}(x_{uv})>\ \tau\\ \text{None},\ \ \ \ \ \text{otherwise}\end{cases}$ (1)
@@ III-C3 Training
For the nuScenes dataset, we first project multiple frames to the current LiDAR frame $\mathbf{d}_{gt}$ to obtain the cumulative $\mathbf{d}_{acc}$. After that, linear interpolation in log space [34] is performed on $\mathbf{d}_{acc}$ to obtain $\mathbf{d}_{int}$. Because of its density, $\mathbf{d}_{gt}$ is directly interpolated without accumulation for the ZJU-4DRadarCam dataset. For supervision, we use $\mathbf{d}_{int}$ to build binary classification labels $\mathbf{y}_{i}\in\left\{0,1\right\}^{H\times W}$, where points with a depth difference less than 0.5m from the Radar point are labeled as positive. After constructing $\mathbf{y}_{i}$, we minimize the binary cross-entropy loss:
@@
where $\Omega\subset\mathbb{R}^{2}$ denotes the spatial image domain of $\mathbf{C}_{i}$, $x\in\Omega$ is a pixel coordinate, and $\mathbf{\hat{y}}_{i}=h_{\theta}(\mathbf{C}_{i},\mathbf{p}_{i})$ is the confidence of correspondence.
@@
Figure 5: (a) Our metric depth estimation over the input image in a cluttered large-scale scenario. (b) Top row shows the ground truth depth $\mathbf{d}_{int}$ and Radar points $\mathbf{P}$ projected into image $\mathbf{I}$. The rest rows from top to bottom depict the depth estimations of [11], [12], and our RadarCam-Depth and the corresponding error maps. Our method demonstrates much higher accuracy and fine details.
@@ III-D2 Training
Ground truth depth $\mathbf{d}_{gt}$ is obtained from projecting 3D LiDAR points. LiDAR depth is further interpolated to get a densified depth $\mathbf{d}_{int}$. During training, we minimize the difference between the estimated metric dense depth $\hat{\mathbf{d}}$ and $\mathbf{d}_{gt}$/$\mathbf{d}_{int}$ with a smoothed L1 penalty:
@@
where $\mathbf{r}_{d}(x)=|\mathbf{d}(x)-\hat{\mathbf{d}}(x)|$, $\lambda_{gt}$ is the weight of $\mathcal{L}_{gt}$, $\Omega_{d}\subset\Omega$ denotes the domains where ground truth has valid depth values. $\beta$ is set to 1 in our practice.
@@ IV-A Datasets
We first evaluate the proposed method on nuScenes benchmark [31]. NuScenes dataset encompasses data collection across 1000 scenes in Boston and Singapore with LiDAR, 3D Radar, camera, and IMU sensors. It comprises around 40000 synchronized Radar-Camera keyframes. We followed the same data splits as [12] with 850 scenes for training and validation and 150 for testing. The test split is officially offered by nuScenes v1.0.
Figure 6: Our data collection routes and the CAD model of our ground robot.
For extensive evaluation, we collected our own dataset, named ZJU-4DRadarCam, using a ground robot (Fig.6) equipped with Oculii’s EAGLE 4D Radar, RealSense D455 camera, and RoboSense M1 LiDAR sensors. Our dataset includes various driving scenarios including urban and wilderness environments. Compared to nuScenes dataset, our ZJU-4DRadarCam offers 4D Radar data with denser measurements, besides, we provide denser LiDAR depth for supervision and evaluation (see Fig.4). Our ZJU-4DRadarCam comprises a total of 33,409 synchronized Radar-Camera keyframes, split into 29312 frames for training and validation, and 4097 frames for testing.

### 19 Mar 2024 v2

@@ Abstract
We present a novel approach for metric dense depth estimation based on the fusion of a single-view image and a sparse, noisy Radar point cloud. The direct fusion of heterogeneous Radar and image data, or their encodings, tends to yield dense depth maps with significant artifacts, blurred boundaries, and suboptimal accuracy. To circumvent this issue, we learn to augment versatile and robust monocular depth prediction with the dense metric scale induced from sparse and noisy Radar data. We propose a Radar-Camera framework for highly accurate and fine-detailed dense depth estimation with four stages, including monocular depth prediction, global scale alignment of monocular depth with sparse Radar points, quasi-dense scale estimation through learning the association between Radar points and image patches, and local scale refinement of dense depth using a scale map learner. Our proposed method significantly outperforms the state-of-the-art Radar-Camera depth estimation methods by reducing the mean absolute error (MAE) of depth estimation by 25.6% and 40.2% on the challenging nuScenes dataset and our self-collected ZJU-4DRadarCam dataset, respectively. Our code and dataset will be released at https://github.com/MMOCKING/RadarCam-Depth.
@@ I Introduction
Perceiving the environment is critically important for autonomous driving, where accurate depth estimation is fundamental for dense reconstruction, 3D detection, and obstacle avoidance. Cameras and range sensors have been widely used for perceiving dense depth. Learned monocular depth (mono-depth) estimation methods based on CNN networks [1, 2, 3, 4, 5] have been prevalent in recent years due to their versatile applicability and plausible accuracy. They benefit from the solid contextual priors from extensive training on diverse datasets. While mono-depth networks excel in estimating up-to-scale depth, they fail to predict the accurate metric scale of depth. This limit arises from the inherent challenge of capturing scale with single-view cameras and the difficulty of learning the diverse scale in complex scenarios.
@@
Figure 1: Top: 3D visualization of the metric depth estimation from our proposed RadarCam-Depth; Middle: Our metric depth estimation overlaid on corresponding error map; Bottom: Error map of Mono-depth after scale-aligned to Radar points. Our depth estimation exhibits exceptional metric accuracy and fine details.
Figure 2: The overall framework of our proposed RadarCam-Depth, comprised with four stages: monocular depth prediction, global alignment of mono-depth with sparse Radar depth, learned quasi-dense scale estimation, and scale map learner for refining local scale. $\mathbf{d}$ and $\mathbf{s}$ denotes the depth and scale, while $\mathbf{z}=1/\mathbf{d}$ is the inverse depth.
@@
The primary contributions of this work are as follows: (i) We introduce the first approach that enhances the highly generalizable, scaleless mono-depth prediction with the dense metric scale intricately inferred from the noisy and sparse Radar data. (ii) We present a novel metric dense depth estimation framework that effectively fuses heterogeneous Radar and camera data. Our framework comprises four stages: mono-depth prediction, global scale alignment of the monocular depth, Radar-Camera quasi-dense scale estimation, and scale map learner for refining the quasi-dense scale locally. (iii) The proposed method is extensively tested on the nuScenes benchmark and our self-collected ZJU-4DRadarCam dataset. It outperforms the state-of-the-art (SOTA) techniques, substantially enhancing Radar-Camera dense depth estimation with high metric accuracy and strong generalizability. (iv) To fertilize future research in robust depth estimation, we will release our code and high-quality ZJU-4DRadarCam dataset, including raw 4D Radar data, RGB images, and meticulously generated ground truth from LiDAR measurements.
@@ II-A Monocular Depth Estimation
Monocular depth estimation is a challenging task due to the inherent scale ambiguity. Many researchers have tried to address this issue by integrating it with optical flow [14], uncertainty estimation [15], semantic segmentation [16], instance segmentation [17] and visual odometry [18]. Although some previous studies [19, 20, 4, 3] have achieved promising results in affine-invariant scaleless depth estimation across diverse datasets, recovering the metric scale remains a significant challenge. Some existing methods rely on inertial data to provide scale. To enhance the generalization, VI-SLAM [21] warps the input image to match the orientation prevailing in the training dataset. CodeVIO [22] proposes a tightly coupled VIO system with optimizable learned dense depth. It jointly estimates VIO poses and optimizes the predicted and encoded dense depth of specific keyframes efficiently. Xie et al. [23] utilize a flow-to-depth layer to refine camera poses and generate depth proposals. They solve a multi-frame triangulation problem to enhance the estimation accuracy. Recently, Wofk et al. [5] introduced a framework for metric dense depth estimation from the VIO sparse depth and monocular depth prediction, which inspires our work. They first globally align the scaleless mono-depth with the metric VIO sparse depth and then learn to refine the dense scale of the globally aligned mono-depth.
@@ II-B Depth Estimation from Radar-Camera Fusion
The fusion of Radar and camera data for metric depth estimation is an active research topic. Lin et al. [8] introduce a two-stage CNN-based pipeline that combines Radar and camera inputs to denoise Radar signals and estimate dense depth. Long et al. [9] propose a Radar-2-Pixel (R2P) network that utilizes radial Doppler velocity and induced optical flow from images to associate Radar points with corresponding pixel regions, enabling the synthesis of full-velocity information. They also achieve image-guided depth completion using Radar and video data [10]. Another approach, DORN [11] proposed by Lo et al., extends Radar points in the elevation dimension and applies deep ordinal regression network-based [24] feature fusion. Unlike other methods, R4dyn [13] creatively incorporates Radar as a weakly supervised signal into a self-supervised framework and employs Radar as an additional input to enhance the robustness. However, their method primarily focuses on vehicle targets and does not fully correlate all Radar points with a larger image area, resulting in lower depth accuracy. Recently, Singh et al. [12] present a method that relies solely on a single image frame and Radar point cloud. Their first-stage network infers the confidence scores of Radar-Pixel correspondence, generating a semi-dense depth map. They further employ a gated fusion network to control the fusion of multi-modal Radar-Camera data and predict the dense depth. However, all the above methods directly encode and concatenate the ambiguous Radar depth and images, confusing the learning pipeline and resulting in suboptimal depth estimation.
@@ III METHODOLOGY
Our goal is to recover the dense depth $\mathbf{\hat{d}}\in\mathbb{R}^{H_{0}\times W_{0}}_{+}$ from a pair of RGB image $\mathbf{I}\in\mathbb{R}^{3\times H_{0}\times W_{0}}$ and Radar point cloud $\mathbf{P}=\left\{\mathbf{p}_{i}|\mathbf{p}_{i}\in\mathbb{R}^{3},i=0,1,2,\cdots,k-1\right\}$ in the image coordinate. $H_{0}$ and $W_{0}$ denote the height and width of the image, respectively. Either 3D or 4D Radar usually has a small field of view with ambiguous sparse data deteriorated by intensive noises. For cross-modal fusion, it is straightforward to project Radar points onto the image plane, generating Radar depth. However, direct fusion of the inherently ambiguous and sparse Radar depth with images, achieved by concatenating their encodings or raw data, can confuse the learning pipeline [25, 26, 27, 12], resulting in aliasing and other undesirable artifacts in the estimated depth. In this paper, we propose to get the scaleless dense depth with existing versatile monocular depth prediction networks, then learn to augment scaleless depth with accurate metric scales from Radar data.
@@ III-A Monocular Depth Prediction
Figure 3: Left: the input image. Middle: Mono-Pred of MiDaS v3.1 [3]. Right: Mono-Pred of DPT-Hybrid [20]. Notably, MiDaS exhibits the ability to differentiate the sky.
@@ III-B Global Alignment
We align the scaleless mono-depth prediction $\mathbf{\hat{d}}_{m}$ with the Radar depth originating from projecting raw Radar points $\mathbf{P}$, by a global scaling factor $\hat{s}_{g}$ and optional offset $\hat{t}_{g}$. The global aligned metric depth is calculated by $\mathbf{\hat{d}}_{ga}=\hat{s}_{g}\cdot\mathbf{\hat{d}}_{m}+\hat{t}_{g}$. Then, it is fed into the subsequent scale map learner (SML). There are many options for performing this global alignment between the projected Radar depth and mono-depth prediction, including: (i) Var: A varying $\hat{s}_{g}$ for individual frame of mono-depth, calculated via root-finding algorithms [29, 30]. (ii) Const: A constant $\hat{s}_{g}$ for all frames of mono-depth prediction, deemed as the mean of scale estimates on the entire training samples. (iii) LS: $\hat{s}_{g}$ and $\hat{t}_{g}$ for individual frames, computed with linear least-squares optimization [19]. (iv) RANSAC: $\hat{s}_{g}$ and $\hat{t}_{g}$ for individual frames, computed with linear least-squares while incorporating RANSAC outlier rejection of the Radar depth. We randomly sample 5 Radar points with valid depth values, estimate $\hat{s}_{g}$ and $\hat{t}_{g}$ with the sampled Radar depth, and adopt the first pair of $\hat{s}_{g}$ and $\hat{t}_{g}$ that yields an inlier ratio over $90\%$. The inlier is the one where the discrepancy between the Radar point depth and aligned mono-depth is under $6$m or the inverse depth discrepancy is under $0.015$.
@@
Figure 4: (a) Top: nuScenes dataset [31] with LiDAR depth $\mathbf{d}_{gt}$ and accumulated LiDAR depth $\mathbf{d}_{acc}$, depth from 3D Radar point cloud $\mathbf{P}$, interpolated LiDAR $\mathbf{d}_{int}$ shown clockwise. The misalignment between LiDAR points and image pixels on this dataset is highlighted with red boxes. Depth from the 3D Radar point cloud is very sparse and non-uniformly distributed. (b) Our ZJU-4DRadarCam dataset with LiDAR depth $\mathbf{d}_{gt}$, interpolated LiDAR depth $\mathbf{d}_{int}$, and depth from 4D Radar point cloud $\mathbf{P}$ shown from top to bottom. Compared to nuScenes, the ZJU-4DRadarCam dataset offers more accurate and denser LiDAR depth and denser 4D Radar depth.
@@ III-C2 Confidence of Cross-Modal Associations
$\hat{\mathbf{d}}_{q}(u,v)=\begin{cases}d(\mathbf{p}_{\mu}),\ \ \ \ \text{if}\ \mathbf{\hat{y}}_{\mu}(x_{uv})>\ \tau\\ \text{None},\ \ \ \ \ \text{otherwise}\end{cases}$ (1)
@@ III-C3 Training
For the nuScenes dataset, we first project multiple frames to the current LiDAR frame $\mathbf{d}_{gt}$ to obtain the cumulative LiDAR depth $\mathbf{d}_{acc}$. After that, linear interpolation in log space [34] is performed on $\mathbf{d}_{acc}$ to obtain $\mathbf{d}_{int}$. Because of its density, $\mathbf{d}_{gt}$ is directly interpolated without accumulation for the ZJU-4DRadarCam dataset. For supervision, we use $\mathbf{d}_{int}$ to build binary classification labels $\mathbf{y}_{i}\in\left\{0,1\right\}^{H\times W}$, where points with a depth difference less than 0.5m from the Radar point are labeled as positive. After constructing $\mathbf{y}_{i}$, we minimize the binary cross-entropy loss:
@@
where $\Omega\subset\mathbb{R}^{2}$ denotes the image region of $\mathbf{C}_{i}$, $x\in\Omega$ is a pixel coordinate, and $\mathbf{\hat{y}}_{i}=h_{\theta}(\mathbf{C}_{i},\mathbf{p}_{i})$ is the confidence of correspondence.
@@
Figure 5: (a) Our metric depth estimation over the input image in a large-scale scenario. (b) Top row shows the ground truth depth $\mathbf{d}_{int}$ and Radar points $\mathbf{P}$ projected into image $\mathbf{I}$. The rest rows from top to bottom depict the depth estimations of [11], [12], and our RadarCam-Depth and the corresponding error maps. Our method demonstrates much higher accuracy and fine details.
@@ III-D2 Training
Ground truth depth $\mathbf{d}_{gt}$ is obtained from projecting 3D LiDAR points. LiDAR depth is further interpolated to get a densified depth $\mathbf{d}_{int}$. During training, we minimize the difference between the estimated metric dense depth $\hat{\mathbf{d}}$ and $\mathbf{d}_{gt}$, $\mathbf{d}_{int}$ with a smoothed L1 penalty:
@@
where $\mathbf{r}_{d}(x)=|\mathbf{d}(x)-\hat{\mathbf{d}}(x)|$, $\lambda_{gt}$ is the weight of $\mathcal{L}_{gt}$, $\Omega_{d}\subset\Omega$ denotes the region where ground truth has valid depth values. $\beta$ is set to 1 in our practice.
@@ IV-A Datasets
IV-A1 NuScenes Dataset
We first evaluate our method on nuScenes benchmark [31]. NuScenes dataset encompasses data collection across 1000 scenes in Boston and Singapore with LiDAR, 3D Radar, camera, and IMU sensors. It comprises around 40000 synchronized Radar-Camera keyframes. We followed the same data splits as [12] with 850 scenes for training and validation, and 150 for testing. The test split is officially offered by nuScenes v1.0.
Figure 6: Our data collection routes and the CAD model of our robot.
IV-A2 ZJU-4DRadarCam Dataset
For extensive evaluation, we collected our own dataset, named ZJU-4DRadarCam, using a ground robot (Fig.6) equipped with Oculii’s EAGLE 4D Radar, RealSense D455 camera, and RoboSense M1 LiDAR sensors. Our dataset consists of various driving scenarios, including urban and wilderness environments. Compared to nuScenes dataset, our ZJU-4DRadarCam offers 4D Radar data with denser measurements. Besides, we provide denser LiDAR depth for supervision and evaluation (see Fig.4). Our ZJU-4DRadarCam comprises a total of 33,409 synchronized Radar-Camera keyframes, split into 29312 frames for training and validation and 4097 frames for testing.